OpenStudy (anonymous):

The sum of two consecutive integers is -25. Find the integers

OpenStudy (anonymous):

x + (x+1) = -25

OpenStudy (anonymous):

you assume that one of the numbers is x, so the other number will be x+1

OpenStudy (anonymous):

Let the first integer be x. And the second integer be x+1. As per the question, x+x+1=-25 Or, 2x=-25-1=-26 So, x=-13 Hence, the consecutive integers are: -12, -13

OpenStudy (anonymous):

x+(x+1)=-25 2x + 1= -25 x =(-25-1)/2 x = -13